if 4 grams of sodium hydroxide is dissolved in 36 grams of water. the solution has the density of 1 gram per ml. find the mole fraction of sodium hydroxide along with molality

Grade:11

1 Answers

Arun
5 years ago
number of moles in 4 g NaOH= wt in gms/molwt=4/40=0.1
number of moles in 36 g H2O=wt in gms/mol wt=36/18=2
total no of moles=0.1+2.0=2.1
mole fraction o​f naoh=0.1/2.1=0.04
mole fraction of h2o=20/21=0.95
 
Density = Mass/ Volume 
Volume = Mass/Density

Volume (solution) = mass of solution/Density of solution
                              = 40 g / 1g ml
                              = 40 g 


Molarity = No . Moles of Solute x 1000 / Volume of the Solution (Into 1000 because Density is in g/ml)
               = .1 x 1000/40 
               = 10/4 
               = 2.5 M